# UOEP: User-Oriented Exploration Policy for Enhancing Long-Term User Experiences in Recommender Systems

### v1

@@ 1 Introduction
Figure 1: Illustration Experiment. We sorted users based on their activity levels (i.e., CTR) and selected the five bottom $\alpha$-quantile ($\alpha\in\{0.2,0.4,0.6,0.8,1.0\}$) of user groups. We then trained two RL algorithms, DDPG and TD3, on these five groups under four varying noise levels. We conducted all experiments with four different seeds and reported and averaged the results. For better presentation, we performed a shift on the values within each group, ensuring that the minimum value within each group becomes 0.1. Adding the value inside the parentheses on the horizontal axis yields the actual return.
@@ 3 UOEP: The Proposed Approach
[…]
@@ 3.3 Facilitating Diversity and Stability for Population
In practice, we will sample a batch of states from the shared replay buffer of the population. We then execute all actors on these sampled states to approximate their respective behavioral embeddings. Once these approximations are obtained, we use a specific metric to quantify the diversity within the actor population. To compare two different actors, we employ a kernel function $\mathcal{K}$ to map their behavioral embeddings into a higher-dimensional feature space. In this space, we compute a kernel matrix $\mathbf{K}$ among the actors within the population, defined as $\mathbf{K}=\mathcal{K}\left(\widehat{\Phi}\left(\pi_{\phi_{i}}\right),\widehat{\Phi}(\pi_{\phi_{j}})\right)_{i,j=1}^{m}$. Geometrically, the determinant of this matrix, $\operatorname{det}(\mathbf{K})$, corresponds to the volume of a parallelepiped defined by the feature maps of the chosen kernel function. A larger volume indicates a higher perceived degree of diversity within the actor population. We optimize population diversity as a regularization term in the population loss function. Specifically, this loss is defined as:
$\mathcal{L}_{\mathrm{div}}(\phi_{1},\ldots,\phi_{m})=-\log\operatorname{det}(\mathbf{K})=-\log\operatorname{det}\left(\mathcal{K}\left(\widehat{\Phi}\left(\pi_{\phi_{i}}\right),\widehat{\Phi}(\pi_{\phi_{j}})\right)_{i,j=1}^{m}\right).$ (10)
@@ 4 Experiments
In this section, we conduct experiments on public and industrial datasets to verify the effectiveness of the proposed UOEP method. We mainly focus on the following questions: Q1. Can UOEP consistently outperform previous state-of-the-art methods (Section 4.2)? Q2. How does UOEP work and how do each of its components contribute (Section 4.3)? Q3. What potential does UOEP have (Section 4.4)?
@@ 4.3 Further Analysis and Ablation Studies
In our method, we employ a distributional critic that allows actors to optimize at different quantile levels of cumulative rewards. To explore the effects of the distributional critic and the role of optimizing at different quantile levels of cumulative rewards, we disabled the distributional critic in UOEP and replaced it with a deterministic critic. In other words, all our actors optimized the expectation of the cumulative return distribution. We provide the recommendation quality of the distributional and deterministic critic in Table 4.3. We can observe a significant decrease in recommendation quality if we use the distributional critic, highlighting the importance of exploration in different user groups.
@@ 4.4 Exploring UOEP’s Potential: Low-Activity User Experience and Fairness
[…]
@@
[…]

### v2

@@ 1 Introduction
Figure 1: Illustration Experiment. We sorted users based on their activity levels (i.e., CTR) and selected the five bottom $\alpha$-quantile ($\alpha\in\{0.2,0.4,0.6,0.8,1.0\}$) of user groups. We then trained two RL algorithms, DDPG and TD3, on these five groups under four varying noise levels. We conducted all experiments with four different seeds and reported the average results. For better presentation, we performed a shift on the values within each group, ensuring that the minimum value within each group becomes 0.1. Adding the value inside the parentheses on the horizontal axis yields the actual return.
@@ 3 UOEP: The Proposed Approach
Figure 2: The proposed approach UOEP. It includes a population of $m$ actors, where $m$ is the population size and each actori outputs an action $a_{i}$ based on the current user state $s$. The action $a_{i}$ along with state $s$ is fed into the distributional critic. Afterward, utilizing both its quantile value $\alpha_{i}$ and the critic’s output $Z(s,a_{i};\cdot)$, actori computes the conditional value at risk (CVaR) measure in order to derive its policy gradients.
@@ 3.3 Facilitating Diversity and Stability for Population
In practice, we will sample a batch of states from the shared replay buffer of the population. We then execute all actors on these sampled states to approximate their respective behavioral embeddings. Once these approximations are obtained, we use a specific metric to quantify the diversity within the actor population. To compare two different actors, we employ a kernel function $\mathcal{K}$ to map their behavioral embeddings into a higher-dimensional feature space. In this space, we choose the squared exponential (SE) kernel, which is defined as follows:
$\mathcal{K}_{\mathrm{SE}}\left(x_{1},x_{2}\right)=\exp\left(-\frac{\left\|x_{1}-x_{2}\right\|^{2}}{2l^{2}}\right),$ (10)
where $l$ is a hyperparameter satisfying $l>0$. Then we compute a kernel matrix $\mathbf{K}$ among the actors within the population, defined as $\mathbf{K}=\mathcal{K}_{\mathrm{SE}}\left(\widehat{\Phi}\left(\pi_{\phi_{i}}\right),\widehat{\Phi}(\pi_{\phi_{j}})\right)_{i,j=1}^{m}$. Geometrically, the determinant of this matrix, $\operatorname{det}(\mathbf{K})$, corresponds to the volume of a parallelepiped defined by the feature maps of the chosen kernel function. A larger volume indicates a higher perceived degree of diversity within the actor population. We optimize population diversity as a regularization term in the population loss function. Specifically, this loss is defined as:
$\mathcal{L}_{\mathrm{div}}(\phi_{1},\ldots,\phi_{m})=-\log\operatorname{det}(\mathbf{K})=-\log\operatorname{det}\left(\mathcal{K}_{\mathrm{SE}}\left(\widehat{\Phi}\left(\pi_{\phi_{i}}\right),\widehat{\Phi}(\pi_{\phi_{j}})\right)_{i,j=1}^{m}\right).$ (11)
@@ 4 Experiments
In this section, we conduct experiments on public and industrial datasets to verify the effectiveness of the proposed UOEP method. We mainly focus on the following questions: Q1. Can UOEP consistently outperform previous state-of-the-art methods (Section 4.2)? Q2. How does UOEP work and how do each of its components contribute (Section 4.3)? Q3. What potential does UOEP have (Section 4.4)? The source code of experiments is shared at https://github.com/lyingCS/UOEP.
@@ 4.3 Further Analysis and Ablation Studies
In our method, we employ a distributional critic that allows actors to optimize at different quantile levels of cumulative rewards. To explore the effects of the distributional critic and the role of optimizing at different quantile levels of cumulative rewards, we disabled the distributional critic in UOEP and replaced it with a deterministic critic. In other words, all our actors optimized the expectation of the cumulative return distribution. We provide the recommendation quality of the distributional and deterministic critic in Table 4.3. We can observe a significant decrease in recommendation quality if we use the deterministic critic, highlighting the importance of exploration in different user groups.
@@ 4.4 Exploring UOEP’s Potential: Low-Activity User Experience and Fairness
CVaR0.3 CVaR0.4 Gini(%) CVaR0.3 CVaR0.4 Gini(%) CVaR0.3 CVaR0.4 Gini(%)
@@
Our actor training leverages the CVaR values from the distributional critic, which focuses on the tail-end of the return distribution. Inspired by recent work that integrates CVaR with collaborative filtering to enhance the experiences of low-activity users (Togashi et al., 2023), we hypothesize that UOEP has similar capabilities for long-term performance, which we validate through empirical analysis. Furthermore, we explore UOEP’s potential to improve the fairness of the long-term performance within the recommender system. We conduct experiments across $3$ datasets with $5$ random seeds. To analyze the model’s performance on low active users, we choose CVaR0.3 and CVaR0.4 as the metrics, which represent the average value of the lower $30$% and $40$% of the Total Reward distribution over the test set, respectively. For assessing fairness, we utilize the Gini coefficient (Wang et al., 2023) as our metric, with lower values indicating greater fairness. Please refer to Appendix F for more details of Gini coefficient. As summarized in Table 4, UOEP outperforms on both the CVaR0.3 and CVaR0.4 metrics across all three datasets. Additionally, the model achieves significant reductions in the Gini coefficient. These results solidly establish UOEP’s effectiveness in serving low-activity users and promoting long-term fairness within the system.